To pay for your​ child's education, you wish to have accumulated ​$14 comma 14,000 at the end of 11 years. To do​ this, you plan to deposit an equal amount into the bank at the end of each year. If the bank is willing to pay 88 percent compounded​ annually, how much must you deposit each year to obtain your​ goal?
Answer #1

FV =14,000
Number of Years =11
If Rate =88%
Then Deposit Per Year =FV/((1+r)^n-1)/r) =14000/((1+88%)^-11)/88%)=11.89

If Rate =8.8%
Then Deposit Per Year =FV/((1+r)^n-1)/r) =14000/((1+8.8%)^-11)/8.8%)=805.85